OpenStudy (anonymous):

A car enters an interstate highway 15 miles north of a city. The car travels north at an average speed of 62.5 mph. Which equation models the car's distance d from the city after traveling for h hours? A) 62.5 = 15h + d B) 15 = h + 62.5 C) d = 62.5h + 15 D) d = 15h + 62.5

OpenStudy (anonymous):

@jalyssa_ar00

OpenStudy (anonymous):

well you know that 62.5 is the rate at which the distance is changing every hour. and you're trying to find d so it's going to be be d equals. something

OpenStudy (anonymous):

15 is your initial condition, since when the car started it was already 15 away

OpenStudy (anonymous):

that's never going to change so it wouldn't make sense for it to give it a variable.

OpenStudy (anonymous):

since h can be anything you'll just leave it as h

OpenStudy (anonymous):

So.. the correct answer would be D ?

OpenStudy (anonymous):

no, because 15 is your starting point and that never changes regardless of h

OpenStudy (anonymous):

Okay so C would be correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks so much. Could you help me with a few more?

OpenStudy (anonymous):

sure, if i can.

OpenStudy (anonymous):

An empty 5-gal water jug weighs 0.75 lb. With 3 c of water inside, the jug weighs 2.25 lb. Which equation models the jug's weight y when it contains x cups of water? A) y = 2x + 0.75 B) y = 0.5x + 0.75 C) y = -x + 3 D) y = -0.167x + 2.25

OpenStudy (anonymous):

well we know that 0.75 is our constant. since the jug by itself always weighs that much.

OpenStudy (anonymous):

lets set up an equation using the givens. \[2.25 = w(3) + 0.75\]

OpenStudy (anonymous):

i used w to signify the rate at which it changes depending on the number of cups.

OpenStudy (anonymous):

alright so when you solve for w what do you get?

OpenStudy (anonymous):

0.75 times 3 equals 2.25 but I'm not sure if that's what you were asking

OpenStudy (anonymous):

Okay so do you understand how i got that equation?

OpenStudy (anonymous):

Yes. Would the answer be B ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Thank you so much. :)

OpenStudy (anonymous):

No problem, I'm glad to help.

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

A concession stand vendor counts the money in the register. After 2 h, it has $82.75. After 6 h, it has $360.75. How much money will be in the register at the end of an 8-hour shift? A) $480.00 B) $489.25 C) $499.75 D) $500.00

OpenStudy (anonymous):

okay so what we're going to have to do here is use some subsitution.

OpenStudy (anonymous):

the equation to model this will be: y = mx + b y is amount m is rate per hour x is hour b is how much you started with

OpenStudy (anonymous):

I messed up. I put A & It was incorrect.

OpenStudy (anonymous):

so using what we're given you get two equations: 82.75 = m(2) + b and 360.75 = m(6) + b

OpenStudy (anonymous):

we can't solve for 2 variables at once so we'll solve for one in terms of the other. So we'll start by converting the first one in terms of b

OpenStudy (anonymous):

Could we skip this one & go to the next one?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

The cost of producing 4 metal units is $204.80. The cost of producing 8 units is $209.60. What is the cost to produce 12 units? A) $212.80 B) $214.00 C) $214.40 D) $216.80

OpenStudy (anonymous):

I think it would be B but I'm not sure...

OpenStudy (anonymous):

this is similar, since we don't know the initial cost. so with the y = mx + b: we get 204.80 = 4m + b and 209.60 = 8m + b

OpenStudy (anonymous):

let's but the first one in terms of b since that's easier. so subtract 4m from both sides: 204.80 - 4m = b

OpenStudy (anonymous):

the 4m on the right side is gone since 4m - 4m is 0 now that we have this equation for b we can plug it into the second equation for b

OpenStudy (anonymous):

so now we get 209.60 = 8m + (204.80 - 4m)

OpenStudy (anonymous):

since the m's have the same variable we subtract them and get 209.60 = 204.80 + 4m

OpenStudy (anonymous):

subtract 204.80 from both sides gets you: 209.60 - 204.80 = 204.80 - 204.80 + 4m 4.80 = 4m

OpenStudy (anonymous):

divide by 4 and m is 1.20

OpenStudy (anonymous):

now we know m = 1.20 so we'll plug it into the other equation 204.80 - 4m = b 204.80 - (4 * 1.20) = b

OpenStudy (anonymous):

204.80 - 4.80 = b b = 200

OpenStudy (anonymous):

so now you have your equation that you can plug a value into: y = 1.20x + 200

OpenStudy (anonymous):

so for 12 units you get y = 1.20(12) + 200

OpenStudy (anonymous):

which becomes y = 14.40 + 200

OpenStudy (anonymous):

So the correct answer would be $214.40 ?

OpenStudy (anonymous):

Yup, you've got it.

OpenStudy (anonymous):

Thank you for explaining that to me :)

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Could you help with 1 more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Look at the scatter plot. Which of these equations is the best representation of the trend line? A) y = 4x + 2 B) y = 1/4x + 2 C) y = -1/4x + 2 D) y = 1/4x - 2

OpenStudy (anonymous):

Do you have a picture of the scatter plot?

OpenStudy (anonymous):

It wont let me get the photo of it.

OpenStudy (anonymous):

But the scatter plot shows the line going through 4 dots with 3 dots above the line & 2 below the line.

OpenStudy (anonymous):

are there number values on it?

OpenStudy (anonymous):

It's just a basic y coordinate, x coordinate graph with the numbers on the left side.

OpenStudy (anonymous):

 something like this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, one sec..

OpenStudy (anonymous):